OpenStudy (anonymous):

((4/5)x)-(3(3/4)(x))= 2(1/2)

OpenStudy (anonymous):

\[4/5x - 3*(3/4x) = 2(1/2)\] Is this the equation?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You can use the equation button below to write an equation clearer than text.

OpenStudy (anonymous):

3*(3/4x) = 9/4x 2*(1/2) = 2/2 = 1 So now you have 4/5x - 9/4x = 1 Now you have to make common denominators. 5 and 4 both go into 20. 4/5x = 16/20x -9/4x = -45/20x 16/20x - 45/20x = -29/20x -29/20x = 1 Now if you multiply both sides by x -29x/20x = 1x The end result is x = -29/20